OpenStudy (anonymous):

The base of a triangle is 4 inches and the area is 20 square inches. What is the height?

OpenStudy (triciaal):

Area of a triangle,A is 1/2 B H where B is the base and H is the height B= 4 and A = 20 20 = 1/2(4)(H) Solve for H

OpenStudy (triciaal):

what did you get?

OpenStudy (anonymous):

Ummmm 2

OpenStudy (triciaal):

how do you let H alone be on one side of the equation?

OpenStudy (triciaal):

20 = 2H

OpenStudy (anonymous):

So the height is 10? idk

OpenStudy (texaschic101):

A = 1/2 BH can be changed to A/(1/2)B = H

OpenStudy (triciaal):

yes check 1/2 (4)(10) = 20

OpenStudy (texaschic101):

10 is correct

OpenStudy (anonymous):

Got it! Thank you both :)

OpenStudy (triciaal):

welcome